40m
QIL
Cryo_Lab
CTN
SUS_Lab
CAML
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m Log |
Not logged in |
 |
|
Fri Oct 7 14:04:48 2011, steve, Update, SUS, He/Ne intensity noise effect on oplevs 7x
|
Fri Oct 7 23:23:05 2011, kiwamu, Update, SUS, SRM oplev was oscillating
|
Wed Oct 19 10:49:35 2011, kiwamu, Update, SUS, SRM oscillation removed
|
Wed Oct 12 08:43:30 2011, steve, Update, SUS, ITMX oplev improved a bit 
|
Thu Oct 13 14:23:09 2011, steve, Update, SUS, ITMX oplev with 3 mm beam on qpd 
|
Mon Oct 17 17:07:30 2011, steve, Update, SUS, ETMX oplev returning beam od 3 mm
|
Wed Oct 19 09:44:10 2011, steve, Update, SUS, ETMX oplev power spectrum
|
Fri May 18 15:25:15 2012, steve, Update, SUS, AR coated lens swapped in at ETMX oplev
|
|
Message ID: 5680
Entry time: Mon Oct 17 17:07:30 2011
In reply to: 5660
Reply to this: 5693
6650
|
Author: |
steve |
Type: |
Update |
Category: |
SUS |
Subject: |
ETMX oplev returning beam od 3 mm |
|
|
ETMX oplev had 6 mm diameter beam on the qpd. I relayed the beam path with 2 lenses to get back 3 mm beam on the qpd
BRC 037 -100 Bi _concave lens and PCX 25 200 VIS do the job. Unfortunately the concave lens has the AR 1064.
|
|
|